Five items was taken from the output of factory. The length and weight of 5 items are given below. Length (meter): 5 6 7 9 12 Weight (KG): 13 15 18 19 20


1
Expert's answer
2022-06-06T13:52:14-0400

In order to compute the regression coefficients, the following table needs to be used:



"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 & X & Y & XY & X^2 & Y^2 \\ \hline
 & 5 & 13 & 65 & 25 & 169 \\
 \hdashline
 & 6 & 15 & 90 & 36 & 225 \\
 \hdashline
 & 7 & 18 & 126 & 49 & 324 \\
 \hdashline
 & 9 & 19 & 171 & 81 & 361 \\
 \hdashline
 & 12 & 20 & 240 & 144 & 400 \\
 \hdashline
Sum= & 39 & 85 & 692 & 335 & 1479 \\
 \hdashline
\end{array}""\bar{X}=\dfrac{1}{n}\sum _{i}X_i=\dfrac{62}{9}=6.89"





"\bar{Y}=\dfrac{1}{n}\sum _{i}Y_i=\dfrac{89}{9}=9.89"





"SS_{XX}=\sum_iX_i^2-\dfrac{1}{n}(\sum _{i}X_i)^2""=335-\dfrac{39^2}{5}=30.8"





"SS_{YY}=\sum_iY_i^2-\dfrac{1}{n}(\sum _{i}Y_i)^2""=1479-\dfrac{85^2}{5}=34"





"SS_{XY}=\sum_iX_iY_i-\dfrac{1}{n}(\sum _{i}X_i)(\sum _{i}Y_i)""=692-\dfrac{39(85)}{5}=29"




"r=\dfrac{SS_{XY}}{\sqrt{SS_{XX}SS_{YY}}}=\dfrac{29}{\sqrt{30.8(34)}}""=0.896155"


Strong positive correlation.